Imperial Jewelers manufactures and sells a gold bracelet for $403.00. The company’s accounting system says that the unit product cost for this bracelet is $264.00 as shown below:

Direct materials $ 141
Direct labor 83
Manufacturing overhead 40
Unit product cost $ 264

The members of a wedding party have approached Imperial Jewelers about buying 22 of these gold bracelets for the discounted price of $363.00 each. The members of the wedding party would like special filigree applied to the bracelets that would require Imperial Jewelers to buy a special tool for $468 and that would increase the direct materials cost per bracelet by $6. The special tool would have no other use once the special order is completed.

To analyze this special order opportunity, Imperial Jewelers has determined that most of its manufacturing overhead is fixed and unaffected by variations in how much jewelry is produced in any given period. However, $7.00 of the overhead is variable with respect to the number of bracelets produced. The company also believes that accepting this order would have no effect on its ability to produce and sell jewelry to other customers. Furthermore, the company could fulfill the wedding party’s order using its existing manufacturing capacity.

Required:

1. What is the financial advantage (disadvantage) of accepting the special order from the wedding party?

2. Should the company accept the special order?
Answer #1

Special order revenue = 22*363 = 7986

Cost of making

= [141+6+83+7]*22 + 468 = 5682

Financial advantage = 7968 - 5682

Financial advantage = 2304

2. Yes